2 black and 3 read pen bought for Rs. 23 then 3 black and 2 read pen bought for Rs. 16 how much paid for 1 black pen?
Answer by rothauserc(4718)

1) 2b +3r = 23
2) 3b +2r = 16
solve second equation for r
2r = 16 -3b
r = 8 -3b/2
substitute for r in first equation
2b +3*(8 -3b/2) = 23
2b +24 -9b/2 = 23
4b/2 -9b/2 = -1
-5b/2 = -1
-5b = -2
b = 2/5
b = .40 rs
now check answer
3*(.40) +2r = 16
2r = 16 - 1.20
r = 7.4
substitute in first equation
2*(.40) +3*(7.40) = 23
.80 + 22.20
answer checks